Question #14278

chances of solving a probability problem indepednently by four students a,b,c,d are 1/2,1/4,1/6 and 1/8.if all of them try to solve the problem .what is the probability that the problem will be solved by at least by one of them

Expert's answer

Probability for all of them being wrong is P=1/2*(1-1/4)*(1-1/6)*(1-1/8)=0.091.
Now probability we're looking for is as follows 1-0.091=0.909
LATEST TUTORIALS
APPROVED BY CLIENTS